# Back Home: A Machine Learning Approach to Seashell Classification and Ecosystem Restoration

### v2

@@ Abstract
In Costa Rica, an average of 5 tons of seashells are extracted from ecosystems annually. Confiscated seashells, cannot be returned to their ecosystems due to the lack of origin recognition. To address this issue, we developed a convolutional neural network (CNN) specifically for seashell identification. We built a dataset from scratch, consisting of approximately 19000 images from the Pacific and Caribbean coasts. Using this dataset, the model achieved a classification accuracy exceeding $85\%$.
The model has been integrated into a user-friendly application, which has classified over 36,000 seashells to date, delivering real-time results within 3 seconds per image. To further enhance the system’s accuracy, an anomaly detection mechanism was incorporated to filter out irrelevant or anomalous inputs, ensuring only valid seashell images are processed.
@@ 1 Introduction
Seashells are essential components of ecosystems, playing a vital role in maintaining ecological balance. As highlighted by Cheng et al. (2023) [3], seashells possess unique structural and functional characteristics, such as high hardness, toughness, corrosion resistance, and bioactivity, which make them crucial for various ecological and industrial applications. Many marine organisms rely on the work of seashells for their survival. Therefore, preserving seashells should be a matter of concern for everyone.
However, in recent years, a significant decline in seashells has been observed due to tourists collecting them from beaches in Costa Rica. Fortunately, these seashells are often confiscated at the Juan Santamaria Airport (Costa Rica’s national airport), but they cannot be returned because their exact ecosystem of origin (Pacific or Caribbean) is unknown.
To address this issue, this project developed a classification model capable of detecting the origin of each species from a single image. A total of 516 species registered in Costa Rica were analyzed, grouped into Pacific and Caribbean categories to perform binary classification. This resulted in a model suitable for returning confiscated seashells at the Juan Santamaria Airport to their respective ecosystems.
@@ 2 Related Work
Several studies have explored the classification of seashells and related marine objects, focusing on aspects such as feature extraction, classification methods, and dataset development. However, these approaches differ from our work, as they primarily emphasize individual feature extraction or taxonomic family classification rather than ecosystem-level insights.
Xue et al. (2021) addressed the challenge of identifying deep-sea debris using deep convolutional neural networks. They introduced the DDI dataset, comprising real deep-sea images, and developed the Shuffle-Xception network. This model incorporated innovative strategies, such as separable convolutions, group convolutions, and channel shuffling, to enhance classification accuracy [22]. Zhang et al. (2019) introduced a pioneering shell dataset containing 7,894 species with over 59,000 images, facilitating research in feature extraction and recognition. They applied traditional machine learning techniques, such as k-NN and random forest, to validate features like color, shape, and texture extracted from the dataset [23]. Yue et al. (2023) proposed ”FLNet”, a convolutional neural network designed to address challenges in shellfish recognition, such as high feature similarity and imbalanced datasets. Their framework included innovative mechanisms for filter pruning and repairing to enhance feature representation, alongside a hybrid loss function tailored for unbalanced datasets. [24]
3 Our Work
The classification model was designed to address the critical challenge of determining the origin of confiscated seashells (Pacific or Caribbean). Due to the specificity of this task, we were required to build a comprehensive dataset from scratch, encompassing nearly every seashell species found in Costa Rica. This effort resulted in the creation of the first seashell database that integrates species from both the Pacific and Caribbean coasts of the country.
By leveraging advancements in convolutional neural network (CNN) architectures and integrating anomaly detection mechanisms, this model ensures high accuracy and reliability for real-world deployment. A user-friendly web application was also developed to facilitate the classification process (see Appendix A). This application enables users to upload images of confiscated seashells and obtain real-time predictions regarding their ecosystem of origin. The app was designed with simplicity in mind, ensuring accessibility for a wide range of users.
This section outlines the data preparation, model selection, and training strategies employed to achieve the desired goal.
3.1 Data Collection and Categorization
The dataset used for this research comprises 19,058 images representing 516 species of seashells, collected over 10 months. From the total of images, 9,553 are from the Caribbean and 9,505 from the Pacific. All of these species were categorized into four distinct groups based on their taxonomy and ecosystem (see Appendix  B)
Category Pacific (Species) Caribbean (Species)
Gastropods 130 149
Bivalves 107 130
Table 1: Comparison of species counts for Gastropods and Bivalves in the Pacific and Caribbean regions.
To construct the dataset, we curated a comprehensive list of species by referencing specialized resources such as Conchology Inc. [11], iNaturalist [10], the Florida Institute of Technology [14], and ConchyliNet [4]. While other seashell datasets exist for image classification [23, 13] they are not tailored to the unique biodiversity of our ecosystems. Therefore, we were required to source specific images for each species, capturing their distinct characteristics and possible variations to ensure accurate representation.
This categorization resulted in a total of 279 species from the Caribbean and 237 species from the Pacific. Each seashell was grouped by family, genus, and species to ensure accurate taxonomic representation. A detailed list of all categorized species is provided in the Appendix for reference. The dataset was meticulously verified to ensure the quality and consistency of the images, incorporating diverse positions, backgrounds, and environmental conditions to enhance the model’s robustness.
Figure 1: Representative Pacific seashells from the dataset.
Figure 2: Representative Caribbean seashells from the dataset.
The dataset was divided into three subsets for training, validation, and testing, with 70% allocated to the training set for model training, 15% to the validation set for hyperparameter tuning and performance evaluation during training, and 15% to the test set for evaluating the final model performance on unseen data. This split ensured a balanced representation of families across all subsets, supporting a rigorous and reliable evaluation of the model’s performance. All images were resized to 224x224 pixels to maintain consistent input dimensions for the model architecture, while preserving key visual features necessary for classification.
To enhance the model’s ability to generalize and handle variations in real-world data, several data augmentation techniques were implemented. These included random rotations (±20 degrees), horizontal flips, and subtle adjustments to brightness and contrast. Additionally, random cropping and zoom operations were applied to simulate different viewing distances and perspectives. These augmentation techniques effectively expanded the training dataset and helped the model become more robust to natural variations in species appearances across different lighting conditions and viewing angles. These categorizations were critical in training the final classification model, as they allowed the system to learn distinctions not only between seashells but also between ecosystems. By leveraging this structured grouping, the model could focus on subtle morphological, color, and texture variations that characterize species from different ecosystems.
3.2 Model Training
The final classification model was built using the ConvNext architecture [12]. We deliberately excluded Vision Transformers (ViT) from our experimentation due to their substantially larger model size and computational requirements, which would have hindered deployment in resource-constrained environments. We experimented with several established architectures including ResNet50 [7], DenseNet121 [8], and MobileNetV2 [18]. Despite extensive training, these models consistently achieved accuracy scores below 81%, even when using identical training configurations (Table 5 - Appendix A).
ConvNext demonstrated superior performance, which we attribute to its 7$\times$7 kernel design that effectively captures global features without relying on computationally expensive attention mechanisms. This choice not only improved accuracy but also reduced the model’s computational overhead, making it more scalable when deployed for many simultaneous users as we did during the classification process.
To train the model we used the ConvNext-Tiny version from PyTorch library with ImageNet-1K weights. The following hyperparameters were used: The learning rate was set to 0.01, and the optimizer used was Stochastic Gradient Descent (SGD) with a momentum of 0.9 and a weight decay of 0.001. The batch size was 16, and the model was trained for 100 epochs. Starting from epoch 65, the learning rate was decayed by a factor of $e^{-0.05}$. Thirty layers were unfrozen during training, allowing for fine-tuning of the deeper layers, while a final dropout layer with a rate of 0.3 was applied to reduce overfitting.
Figure 3: Classification process utilizing a modified ConvNext Tiny architecture
3.3 Anomaly Detection System
Anomaly systems have been broadly studied over recent years, especially for Large Language Models (LLMs). They have shown promising results in filtering out inputs that differ substantially from labeled datasets used for training [21, 20, 19].
Image analysis in real-world scenarios often requires systems that can identify when an input cannot be reliably assigned to any of the known classes, which is crucial in classification or detection tasks deployed in practical settings. Deep learning has emerged as a key enabler for such anomaly detection systems, such as Autoencoders and Generative Adversarial Networks to learn latent representations of normal data. By modeling what is typical, these methods can highlight inputs that fall outside the learned distribution. [17, 1]
Other approaches employ neural networks to estimate elements such as depth, perform feature extraction, or localize specific regions within an image, thereby providing a representation of the anomalous portion [2, 15]. Also, certain approaches emphasize feature extraction and are thus more aligned with our goal: to obtain a robust image representation and subsequently determine whether an input is anomalous relative to a trained reference dataset. [6, 16, 5]
@@ 3.3 Anomaly Detection System
Our approach focuses on using a vector representation of each training image and storing these vectors in a vector database. This enables efficient comparison against images provided by volunteers, resulting in a more accurate classification based on their similarities. This method ensures that only images containing the seashell as the primary element of interest are selected, reducing the impact of extraneous noise that could adversely affect the classification. Such an approach is particularly beneficial for a web application utilized by a wide range of users, many of whom may lack professional or scientific photography skills.
This system identifies images that do not correspond to seashells, preventing incorrect classifications during the real-time volunteering. The process involves creating image embeddings using a convolutional neural network (SqueezeNet1.0) [9]. Each image was transformed into a vector of 1,000 scalar values, representing a compact and discriminative feature set.
The embeddings were stored in a vector database to facilitate efficient similarity comparisons. When a new image is provided, its embedding is generated and compared against the stored embeddings using cosine similarity. The anomaly detection system operates as follows:
  1. 1.
    Compute the cosine similarity between the new image embedding and the top $k$ closest embeddings in the database.
  2. 2.
    Average the similarity scores. If the resulting score is below a threshold ($\lambda$), the image is classified as an anomaly.
Category Images Below Threshold*

### v3

# Back Home: A Computer Vision Solution to Seashell Identification for Ecological Restoration

@@ Abstract
Illegal souvenir collection strips an estimated five tonnes of seashells from Costa Rica’s beaches each year, yet once these specimens are seized their coastal origin—Pacific or Caribbean—cannot be verified easily because the lack of information, preventing their return when are confiscated by the local authorities. To solve this issue, we introduce BackHome19K, the first large-scale image dataset annotated with coast-level labels, and propose a lightweight pipeline that infers provenance in real time on a mobile-grade CPU. A similar PaDiM-inspired anomaly filter pre-screens uploads, increasing robustness to user-generated noise. On a held-out test set the classifier attains 86.3 % balanced accuracy, while the filter rejects 93 % of 180 out-of-domain objects with zero false negatives. Deployed as a web application, the system has already processed 70 000 shells for wildlife officers in under three seconds per image, enabling confiscated specimens to be safely repatriated to their native ecosystems. The dataset is available at FIFCO/BackHome19K.
@@ 1 Introduction
Figure 1: Web interface displaying the final output of our pipeline: for each uploaded seashell image the system returns its predicted coastal provenance with a confidence score
Figure 2: Overview of the proposed two-stage pipeline for seashell origin classification. The system first applies anomaly detection to identify and filter out abnormal specimens, then employs a classification network to predict the geographic origin of authentic seashells. The integrated approach ensures robust performance by eliminating outliers before classification.
Seashells are a keystone element of coastal ecosystems. Beyond providing habitat and shelter, their biological structures regulate sediment chemistry, buffer pH, and contribute to nutrient cycling. Recent materials-science studies [6, 22] have further shown that molluscan shells possess exceptional hardness, fracture toughness, corrosion resistance, and bio activity. Preserving this natural capital is therefore critical for biodiversity, shoreline stability, and sustainable innovation.
Despite their importance, seashell populations along Costa Rica’s beaches have sharply declined owing to unregulated souvenir collection by tourists. Customs officers at Juan Santamaría International Airport routinely confiscate these shells, yet repatriation is impossible because the specimens’ provenance—Pacific or Caribbean coast—cannot be determined post-hoc. Returning them to the wrong habitat risks spreading invasive parasites, disrupting local gene pools, and skewing ecological surveys [24, 30]. Consequently, thousands of shells remain in storage each year instead of being restored to their native ecosystems [41, 1, 42]
This conservation bottleneck presents a classic computer vision challenge: fine-grained classification where visually similar species from different ecosystems must be distinguished based on subtle morphological differences [17, 3]. Traditional taxonomic identification focuses on species-level classification, but conservation applications require ecosystem-level provenance determination that demands specialized datasets and architectures optimized for geographic rather than purely biological distinctions.
The challenge extends beyond traditional fine-grained classification because Pacific and Caribbean shells often share nearly identical morphological features—similar color palettes, growth ring patterns, and overall silhouettes—yet originate from fundamentally different marine ecosystems [14, 8]. This requires our model to learn extremely subtle distinguishing characteristics: minute texture variations, slight hue shifts, and microscopic geometric deviations that reflect different environmental conditions, water chemistry, and ecological pressures.
We tackle this logistical and conservation bottleneck through an image-based classification model. Specifically, we formulate the task as a binary recognition problem—Pacific versus Caribbean—and introduce a large-scale dataset comprising $\sim$19 000 high-resolution photographs containing 516 mollusk species presented in each location. Building on this resource, we design a lightweight convolutional neural network (CNN) tailored to the fine-grained visual cues that distinguish shells from the two coasts.
Our contributions are threefold:
  1. 1.
    Dataset. We release BackHome19K, the first annotated image corpus of Costa Rican seashells with coast-level labels, captured in situ under natural lighting conditions across 516 species.
  2. 2.
    Model. We propose a compact CNN architecture based on ConvNeXt-Tiny that balances classification accuracy with the low-latency constraints of field deployment.
  3. 3.
    System. We integrate the model into a production-ready mobile and web application that incorporates anomaly detection for robust real-world performance, enabling wildlife authorities to triage confiscated shells in real time and streamline the path from seizure to ecosystem restoration.
@@ 2 Related Work
Our research builds upon two primary areas: ecological classification models and image embeddings for real-time anomaly detection systems. The former involves techniques for identifying species or ecosystems based on visual or environmental data, while the latter focuses on methods for detecting outliers or non-conforming inputs to ensure reliability and robustness in real-world deployments.
2.1 Architectures
Convolutional Neural Networks (CNNs) have been central to image classification since the breakthrough by [23]. Efficient variants like ResNet [15], MobileNetV2 [38], and DenseNet [16] improved performance using residual connections, depthwise separable convolutions, and dense connectivity.
ConvNeXt [28] modernized CNNs by integrating transformer-inspired elements (large kernels, LayerNorm, GELU), narrowing the performance gap with transformers.
Vision Transformers (ViTs) [11], and later Swin [27] and DeiT [44], replaced convolutions with attention-based mechanisms, achieving state-of-the-art results by modeling global patch relationships. This shift from CNNs to transformer-based models continues to improve recognition tasks across domains.
2.2 Ecology Classification Models
Initial research framed marine shell recognition as a fine-grained classification task. [46] introduced the DDI dataset and proposed Shuffle-Xception, enhancing accuracy on cluttered seabed scenes [47]. [49] compiled a large benchmark (7,894 species, 59,000 images), evaluating traditional descriptors with classical classifiers.
To address class imbalance and visual similarity, [48] proposed FLNet, a CNN with filter pruning and a hybrid loss tailored to skewed distributions. More recently, [12] applied transformers for fish recognition, leveraging transfer learning for high accuracy.
Despite progress, most approaches emphasize taxonomy and overlook ecological context—such as geographic provenance—which is crucial for conservation and real-world deployment.
2.3 Image Embeddings and Anomaly Detection
Deep learning enables encoding images as high-dimensional embeddings that capture semantic content, supporting tasks like retrieval, classification, and anomaly detection. These embeddings cluster semantically similar inputs in feature space, offering more meaningful comparisons than raw pixels. Self-supervised methods have largely superseded supervised CNNs by learning from unlabeled data.
SimCLR [5] leveraged contrastive learning to encode image semantics via augmented views. CLIP [33] extended this to multimodal embeddings, enabling zero-shot transfer across vision-language tasks. DINO [4] showed that vision transformers can learn structural and semantic features without labels.
For scalable retrieval, embeddings are indexed using approximate nearest neighbor (ANN) algorithms like FAISS, enabling efficient search in large datasets [34].
Embeddings also underpin anomaly detection—key to identifying out-of-distribution (OOD) inputs. Similar techniques are used in NLP to detect hallucinations in LLMs [43, 40, 39]. Visual anomaly detection methods fall into three categories: (i) reconstruction-based (autoencoders, VAEs [37, 2]), (ii) probabilistic modeling (PaDiM [9], CS-Flow [36]), and (iii) confidence-based methods like ODIN [25].
While not central to our system, anomaly detection modules improve robustness against unexpected user-generated inputs in field conditions.
3 Method
This work makes three key contributions. First, we curate and release the first large-scale seashell image corpus, encompassing photographs from seashell species collected from both Costa Rica’s Pacific and Caribbean coasts—an essential resource for future marine-biodiversity studies. Second, we trained a lightweight ConvNeXt-Tiny–based classifier that distinguishes between the two ecosystems even for visually similar species from the same family that inhabit different shores. Third, we wrap the model in a production-ready web service that delivers predictions in under three seconds per image and incorporates an integrated anomaly detector, ensuring corrupted or out-of-distribution inputs are flagged before they reach the classifier.
3.1 Dataset Construction
We introduce BackHome19K, the first comprehensive seashell dataset specifically designed for ecosystem-level classification. Our dataset addresses a critical gap in existing marine datasets [49, 31], which focus on taxonomic classification rather than bio geographic origin inference.
Figure 3: Representative examples of bivalves and gastropods collected from Caribbean coasts.
Because Pacific and Caribbean shells often share near-identical color palettes, growth rings, and overall silhouettes, a small dataset would leave the model unable to tease apart these subtle cues. We therefore assembled a broad and balanced collection spanning 516 species based on a list of seashells provided by the Biology School from the Universidad de Costa Rica.
Our dataset encompasses a diverse collection of marine mollusks from Pacific and Caribbean waters, representing major taxonomic groups across multiple families. Among the gastropods, we include specimens from families such as Acmaeidae (limpets), Buccinidae (whelks), Bullidae (bubble shells), Calliostomatidae (top shells), Cassidae (helmet shells), Cypraeidae (cowries), and Fasciolariidae (spindle shells), among numerous others. The bivalve collection features representatives from Arcidae (ark shells), Ostreidae (oysters), Pectinidae (scallops), Tellinidae (tellins), Semelidae (semele clams), and Veneridae (venus clams), providing comprehensive coverage of both filter-feeding and burrowing species across varied marine habitats.
Figure 4: Representative examples of bivalves and gastropods collected from Pacific coasts.
We systematically cataloged the species endemic to Pacific and Caribbean coasts through consultation with marine biology experts and cross-referencing multiple authoritative databases [20, 19, 32, 7]. Species selection criteria included: (1) confirmed presence in target ecosystems, (2) sufficient morphological distinctiveness for visual classification, and (3) availability of high-quality reference imagery.
Figure 5: Seashell family clusters based on the mean feature vectors for each specie using embedding representations from all the images in the dataset using t-SNE [45] and DBSCAN [13]
The dataset was assembled by three researchers under the guidance of Dr. Yolanda Camacho García, who provided the species list. Using her taxonomic expertise, the lead researcher defined a labeling protocol to ensure consistency within classes and clear separation between them. Images were chosen based on key morphological traits—texture, color, shape, and distinctive external features—and organized into class‑specific folders grouped by ecosystem (Pacific vs. Caribbean). In Figure 5, we then computed each family’s mean embedding and applied t‑SNE to visualize clusters of families with similar color, shape, texture, and size.
Figure 6: Representative specimens from our dataset showing morphological diversity across families. Top row (left to right): Pectinidae (scallop with characteristic radial ribs), Cardiidae (cockle with prominent radial sculpture), Cardiidae (heart cockle with distinctive coloration), Veneridae (venus clam with concentric growth lines). Middle row: Fasciolariidae (tulip shell with elongated spire), Cypraeidae (cowrie with glossy, oval shell), Turbinidae (turban shell with nacreous interior visible). Bottom row: Mitridae (mitre shell with geometric pattern), Turritellidae (tower shell with high spire), Tellinidae (tellin with smooth, elongated form), Buccinidae (whelk with pointed spire), Fasciolariidae (spindle shell with fusiform shape).
Over 10 months, we assembled 19,051 images. Each of them underwent rigorous quality control including: taxonomic verification, removal of synthetic/composite images and standardization of lighting conditions To ensure morphological diversity, we collected 30–40 pictures per species showing different growth stages, orientations, and preservation states.
The dataset was divided into Pacific and Caribbean subsets for the specific purpose of training the model introduced in the next section. However, the images are also organized by family, genus, and species, providing broader opportunities for future research. This structure enables studies focused on inter-species similarities and supports classification tasks based on taxonomic attributes rather than geographic origin.
Table 1: Dataset composition and characteristics for Pacific and Caribbean divided into Gastropods and Bivalves
Metric Pacific Caribbean
Total Species 237 279
Total Images 9,505 9,553
Gastropod Species 130 149
Bivalve Species 107 130
Avg. Images/Species 40.1 34.2
3.2 Model Architecture
We adopt ConvNext-Tiny [28] as our classification backbone, initialized with ImageNet-1K pre-trained weights. Our architectural choice addresses two critical deployment constraints: computational efficiency for real-time inference and model size limitations for mobile deployment scenarios.
ConvNext modernizes traditional CNN design by incorporating key innovations from Vision Transformers while maintaining computational efficiency. The architecture employs large 7$\times$7 convolution kernels in early stages, enabling effective global feature capture without the quadratic complexity of self-attention mechanisms. This design proves particularly advantageous for fine-grained classification tasks where subtle morphological differences must be detected across varying spatial scales. We deliberately exclude Vision Transformers from consideration due to model size, ViT-Base requires 86M parameters vs. ConvNext-Tiny’s 28M, creating storage constraints for field deployment; second, self-attention scales quadratically with input resolution, limiting real-time performance on resource-constrained devices.
This architecture already encodes a rich hierarchy of visual cues, from global shape and size in its early layers to increasingly fine color bands and micro-textures in its final block. By freezing the stem and first three stages we lock in the generic edge, contour, and chromatic primitives that distinguish broad shell geometries, while selectively unfreezing only the last block to relearn the millimeter-scale ridges, growth rings, and pigment speckles that separate near-identical families from Costa Rica’s Pacific and Caribbean coasts.
3.3 Anomaly Detection
@@
Real-world deployment necessitates filtering non-seashell inputs that could degrade classification performance. Therefore, our approach focuses on using a vector representation of each training image and storing these vectors in a vector database. This enables efficient comparison against images provided by volunteers, resulting in a more accurate classification based on their similarities. This method ensures that only images containing the seashell as the primary element of interest are selected, reducing the impact of extraneous noise that could adversely affect the classification. Such an approach is particularly beneficial for a web application utilized by a wide range of users, many of whom may lack professional or scientific photography skills.
Our anomaly–detection module adopts the embedding logic of PaDiM [9] while using a more lightweight backbone. Concretely, each image is forwarded through SqueezeNet 1.0 [18]; in line with PaDiM we take the global-average-pooled activations of the final convolutional layer as the feature vector, yielding a 1 000-dimensional descriptor.
The choice of SqueezeNet is driven by its compactness and representational strength. With only 1.2$\approx$ 1.2≈ 1.2 M parameters—roughly 20×20$\times$20 × fewer than canonical backbones such as ResNet-50—it fits comfortably on resource-constrained devices. In addition, its Fire modules (1 $\times$ 1 “squeeze” followed by 1 $\times$ 1 / 3 $\times$ 3 “expand” filters) and delayed down-sampling keep high-resolution feature maps deep into the network, enabling it to capture the fine local textures that reveal subtle, pixel-level anomalies.
Having obtained a 1 000-dimensional embedding $\mathbf{e}_{q}$ for each query image, the system then measures its affinity to the dataset by computing the mean cosine similarity to its $k$ nearest neighbors:
$S\;=\;\frac{1}{k}\,\sum_{i=1}^{k}\frac{\langle\mathbf{e}_{q},\mathbf{e}_{i}\rangle}{\lVert\mathbf{e}_{q}\rVert\,\lVert\mathbf{e}_{i}\rVert}.$
This method ensures that only valid seashell images are processed by the classification system, improving overall accuracy and reliability. The threshold $\lambda$ was determined empirically by analyzing the distribution of similarity scores between known seashell images in our dataset.
Figure 7: Feature extraction and embedding generation from input images using the SqueezeNet architecture
4 Experiments
In the following section we indicate the procedure for the model training and hyperparameter selection.
4.1 Dataset Setting
The dataset was divided into three subsets for training, validation, and testing, with 70% allocated to the training set for model training, 15% to the validation set for hyperparameter tuning and performance evaluation during training, and 15% to the test set for evaluating the final model performance on unseen data. This split ensured a balanced representation of families across all subsets, supporting a rigorous and reliable evaluation of the model’s performance. The decision to retain almost 2,900 pictures for testing was made because of the need to ensure that the model would perform correctly on completely new seashells, as it was required to be used in a real-time framework with completely new seashells, ensuring that our dataset distribution could cover those undiscovered seashells. In addition, all images were resized to 224x224 pixels to maintain consistent input dimensions with the model architecture.
[…]
4.2 Implementation Details and Hyperparameter Analysis
All experiments were implemented in PyTorch and conducted on a single NVIDIA A100 GPU. We performed extensive hyperparameter optimization across seven distinct configurations, systematically evaluating the impact of unfrozen layer counts, training duration, and learning rate scheduling on model performance.
We evaluated three optimizers (AdamW [29], Adam [21], and Stochastic Gradient Descent (SGD) [35]) with initial learning rates spanning $1\times 10^{-4}$ to $1\times 10^{-2}$, comparing two scheduling strategies: cosine annealing and ReduceLROnPlateau. Our final configuration employs SGD with an initial learning rate of 0.001 and cosine annealing, which demonstrated superior convergence characteristics.
The architectural decisions were guided by empirical findings:
  • Unfreezing the final 30 layers of ConvNext-Tiny provided optimal feature adaptation without overfitting
  • Strategically placed dropout layers mitigated overfitting on underrepresented seashell families
  • The chosen configuration particularly benefits fine-grained classification, where subtle inter-class differences demand careful feature extraction
This systematic exploration establishes a robust baseline for seashell classification while providing insights into transfer learning optimization for fine-grained visual tasks.
5 Results
5.1 Classification Performance
As shown in table 3, the model achieves balanced performance across both ecosystems, with slightly higher accuracy for Caribbean specimens (87.10% vs. 85.43%). This difference may reflect the larger number of Caribbean species in our training set (279 vs. 237 Pacific species) .
Table 2: Top-1 test accuracy (mean ± std across ten experiments) of four backbone CNNs. ConvNeXt-Tiny outperforms the next-best DenseNet121 by roughly 6 percentage points.
Architecture Test Accuracy (%)
ResNet50 78.3 $\pm$ 0.4
DenseNet121 80.2 $\pm$ 0.3
MobileNetV2 79.3 $\pm$ 0.5
ConvNext-Tiny 86.28 $\pm$ 0.5
ConvNeXt-Tiny’s superior performance stems from its modern architectural design, which effectively preserves high-frequency visual details crucial for fine-grained classification tasks. The network’s large 7×7 convolution kernels in early stages enable comprehensive feature capture, while its hierarchical structure maintains discriminative micro-textures and subtle morphological variations that distinguish shells from different ecosystems.
In contrast, traditional architectures like ResNet50 [15] and MobileNetV2 [38] struggle to preserve these fine-grained details due to their aggressive down-sampling strategies and smaller receptive fields. This limitation becomes particularly problematic when classifying morphologically similar seashells from the same taxonomic family but different geographic origins, where the distinguishing features may be limited to minute surface textures.
Table 3: Classification performance on BackHome19K test set (2,858 images) showing balanced accuracy across Pacific and Caribbean ecosystems
Ecosystem Accuracy (%) Precision (%) Recall (%) F1 (%)
Pacific 85.43 88.70 85.43 87.00
Caribbean 87.10 83.45 87.10 85.29
Overall 86.28 86.23 86.13 86.17
The results demonstrate that the model performs consistently across both ecosystems, achieving a balanced accuracy of 85.43% on Pacific samples and 87.10% on Caribbean samples. This indicates that the system correctly classifies nearly 9 out of 10 seashells into their respective native ecosystems, highlighting its robustness and generalization capabilities across diverse marine environments.
5.2 Anomaly Detection Performance
Our anomaly detection system successfully filters non-seashell inputs while preserving legitimate specimens. Testing on 200 images across 20 object categories (10 images each, except 40 seashell images), the system achieved 100% recall on seashell images while correctly identifying 90.5% of non-seashell objects as anomalies using the empirically determined threshold $\lambda=0.955$ and $k=5$
To probe the robustness of our similarity–based filter, we curated an out-of-domain corpus from different standard vision benchmarks such as COCO [26], ImageNet-1k [10], and Places365 [50]. For each of 18 everyday object categories, we sampled ten high-quality photos, yielding 180 non-seashell test images. We also added forty previously unseen in-domain seashell pictures for a recall check.
Figure 8: Anomaly detection interface displaying rejection of non-seashell objects using similarity-based filtering with threshold $\lambda=0.955$, achieving 93% precision on out-of-domain images
A fixed threshold of 0.955 rejected 168 of the 180 off-domain images (93 % true-positive rate). Perfect rejection was achieved for visually distinctive classes such as cats, cars, dogs, trucks, and backgrounds (10 / 10 each). More heterogeneous textures—notably reptiles, people, and frogs—accounted for the remaining misses, yet still exceeded a 60 % rejection rate. Crucially, all forty seashell controls scored above the threshold, producing zero false negatives and confirming that the detector preserves recall on in-domain data while aggressively filtering other objects.
6 Web Application
[…]
The entire stack, React assets, FastAPI render and predict containers, is Dockerized and deployed to Cloud Run, leveraging scale to zero, sub second cold starts, and horizontal auto scaling. This architecture ensures sub 3 seconds end to end response times under heavy concurrency while minimizing idle costs.
Figure 9: End-to-end technical pipeline for real-time shell classification system with sub-3-second response times via distributed FastAPI services and Cloud Run deployment
[…]
7 Ablations
To evaluate the performance of the ConvNext-Tiny architecture and our decission to unfreeze a set of layers given the efficiency and considering the benefits of initial ones, we ran a set of experiments using the ConvNext-Tiny architecture introduced above with a set of same hyperparameters to evaluate the changes. We trained these models with SGD [35] and with a learning rate of $1\times 10^{-2}$ and just differing the learning rate epoch scheduler and the number of epochs.
Table 4: Ablation Study: Model Performance with Different Configurations to observe unfrozen layers importance
Unfrozen Epochs Scheduling Accuracy
0 50 25 83.24
0 50 25 84.11
10 100 50 80.15
10 150 75 84.56
10 250 125 84.16
13 150 75 84.16
30 100 68 86.28
Based on Table 4, we observe that unfreezing 30 layers yields the best performance with 86.28% accuracy. Notably, this configuration required only 100 training epochs compared to the 13-layer unfrozen experiment, which trained for 150 epochs yet achieved 2.12% lower accuracy (84.16%). The accuracy is presented on the test dataset.
The results demonstrate a clear relationship between the number of unfrozen layers and model performance. Experiments with 10 unfrozen layers consistently underperformed compared to the 30-layer configuration, with accuracies ranging from 80.15% to 84.56%. While the baseline frozen models (0 unfrozen layers) achieved reasonable performance (83.24%–84.11%), they were still outperformed by the optimal unfrozen configuration.
It is important to note that the varying number of epochs across experiments was determined based on validation accuracy convergence patterns observed during training using an early stop function. Each configuration was trained until the validation accuracy plateaued, indicating that additional epochs would incur unnecessary computational costs without performance gains.
8 Limitations
While the classification system demonstrated strong performance, several limitations were observed. The model occasionally failed under suboptimal lighting conditions or unconventional viewing angles/distances, significantly affecting extracted visual features and reducing accuracy. Additionally, some Caribbean seashells lack sufficient training images in research databases, creating classification difficulties.
Although the anomaly detection mechanism filtered many irrelevant inputs, it was not infallible. Non-seashell objects with similar textures (coral fragments, rocks, marine debris) were sometimes misclassified as seashells, indicating the need for better fine-grained visual distinction.
Figure 10: Example of seashells that were misclassified into Caribbean due to challenging conditions such as viewpoint, depth, lighting, and lack of visible texture or detail. From left to right: Pectinidae, Arcidae, Cypraea (×2), and a small Cardiidae shell.
The current dataset includes 516 species based on available scientific knowledge, though this initial version may not cover all existing species. Future updates will address these gaps as new discoveries emerge.
9 Conclusion
This research demonstrates that modern computer vision can effectively support conservation efforts through practical, deployable systems. Our BackHome19K dataset and classification pipeline achieve 86.3% accuracy in determining seashell ecosystem provenance, enabling the automated return of confiscated specimens to their native habitats.
This work establishes a framework for ecosystem-level classification that could extend to other marine organisms and geographic regions. The integration of anomaly detection with fine-grained classification proves essential for real-world deployment, where user-generated content varies significantly in quality and relevance.
The public release of BackHome19K will enable comparative studies and methodological improvements across the computer vision for ecology community.
Acknowledgments
We would like to thank FIFCO and the advertising agency Joystick for the funding and support provided throughout the duration of this project. We are especially grateful to Yolanda Camacho, PhD for her invaluable work and dedication in compiling the seashell lists. We also express our appreciation to the Sistema Nacional de Áreas de Conservación (SINAC), the Universidad de Costa Rica (UCR), and AERIS for their collaboration and assistance throughout this process. Finally, we extend our thanks to Vivian Mayorga from Joystick for her support in designing the web application and visual elements of this paper.
10 Supplementary Material
10.1 Anomaly Detection Algorithm
Category Images Below Threshold